# Prioritized Information Bottleneck Theoretic Framework with Distributed Online Learning for Edge Video Analytics

### v1

@@ Abstract
Collaborative perception systems leverage multiple edge devices, such surveillance cameras or autonomous cars, to enhance sensing quality and eliminate blind spots. Despite their advantages, challenges such as limited channel capacity and data redundancy impede their effectiveness. To address these issues, we introduce the Prioritized Information Bottleneck (PIB) framework for edge video analytics. This framework prioritizes the shared data based on the signal-to-noise ratio (SNR) and camera coverage of the region of interest (RoI), reducing spatial-temporal data redundancy to transmit only essential information. This strategy avoids the need for video reconstruction at edge servers and maintains low latency. It leverages a deterministic information bottleneck method to extract compact, relevant features, balancing informativeness and communication costs. For high-dimensional data, we apply variational approximations for practical optimization. To reduce communication costs in fluctuating connections, we propose a gate mechanism based on distributed online learning (DOL) to filter out less informative messages and efficiently select edge servers. Moreover, we establish the asymptotic optimality of DOL by proving the sublinearity of their regrets. Compared to five coding methods for image and video compression, PIB improves mean object detection accuracy (MODA) while reducing 17.8% and reduces communication costs by 82.80% under poor channel conditions.
@@ I-A Background
Video analytics is rapidly transforming various sectors such as urban planning, retail analysis, and autonomous navigation by converting visual data streams into useful insights[2]. A large number of video cameras produce vast amounts of video data continuously and often require real-time video stream[3]. Numerous developing applications such as remote patient care[4], video games[5], and virtual and augmented reality depend on the efficient analysis of video data with minimal delay[6].
@@
The increasing number of smart devices requires a computational paradigm shift towards edge computing. This approach involves processing· data closer to its source, resulting in several benefits compared to traditional cloud-based paradigms, particularly reduced latency and bandwidth costs. Even a delay as short as second can lead to disastrous consequences. For example, interactive applications such as online gaming and video conferencing require latencies below 100 ms to ensure real-time feedback and seamless user experience[7]. Similarly, VR/AR applications demand extremely low latencies, often less than 20 ms, to prevent motion sickness and maintain a high-quality immersive experience[8]. Utilizing remote cloud services for data processing can result in significant latency increases, often exceeding 100 ms[9]. Moreover, the importance of privacy, particularly in regions with strict data protection laws such as the General Data Protection Regulation (GDPR), makes edge computing even more attractive[10]. According to the Ponemon Institute, 60% of companies express apprehension toward cloud security and decide to manage their own data onsites in order to mitigate potential risks[11].
@@
However, the integration of edge devices into video analytics also brings in many significant challenges[12]. The computational demands of deep neural network (DNN) models, such as GoogLeNet[13], which requires about 1.5 billion operations per image classification, place a substantial burden on the limited processing capacities of edge devices[14]. Additionally, the outputs from high-resolution cameras increase the communication load. For example, a 4K video stream requires up to 18 Gbps of bandwidth to transmit raw video data, potentially overwhelming wireless networks[15]. Therefore, we need to explore efficient video coding for compressing streamed videos. As shown in Fig. 1(a), the traditional compression is to reconstruct streaming frame through efficient entropy coding and motion prediction. However, there are still extensive less informative data being processed, wasting communication bandwidth. For instance, if the tasks involve human recognition or positioning, the reconstructed background of each frame might not be useful for the application.
@@
In this paper, we aim to refine multi-camera video analytics by developing a strategy to prioritize wireless video transmissions. Our proposed Prioritized Information Bottleneck (PIB) strategy attempts to effectively leverage SNR and RoI to selectively transmit data features, significantly reducing computational load and data transmissions. Our method can decrease data transmissions by up to 82.80%, while simultaneously enhancing the mean object detection accuracy (MODA) compared to current state-of-the-art techniques. This approach not only compresses data but also intelligently selects data for processing to ensure only relevant information is transmitted, thus mitigating noise-induced inaccuracies in collaborative sensing scenarios. This innovation sets a new benchmark for efficient and accurate video analytics at the edge.
@@ I-C Our Contributions
  • Our extensive experimental evaluations demonstrate that PIB significantly enhances mean object detection accuracy (MODA) and reduces communication costs. Compared to TOCOM-TEM, JPEG, H.264, H.265, and AV1, PIB improves MODA by 17.8% while reducing communication costs by 82.80% under poor channel conditions. Additionally, our method can reduce the standard deviation of streaming packet sizes by up to 9.43%, while simultaneously maintaining higher MODA, ensuring better transmission robustness under poor channel conditions.
@@ II System Model
As illustrated in Fig. 2, our system comprises a set of edge cameras, denoted as $\mathcal{K}=\{1,2,\ldots,K\}$, and a set of edge servers, denoted as $\mathcal{S}=\{1,2,\ldots,S\}$. Each camera has a specific Field of View (FoV), $\text{FoV}_{k}$, covering a subset of the total monitored area. In a high-density pedestrian environment, our goal is to facilitate collaborative perception for pedestrian occupancy prediction, under the constraints of limited channel capacity due to poor channel conditions. Moreover, the aggregate server is the central fusion edge server selected to minimize overall system delay. Other edge servers relay their data to this server for inference.
@@ II-B Priority Weight Formulation
Dynamic priority weighting is essential for optimizing network resource allocation, as various data sources require different levels of attention. Inspired by our previous work[18], we employ a dual-layer Multilayer Perceptron (MLP) [^1] The MLP is trained in a supervised learning manner, where the input features are the normalized delay ${d}_{\text{norm},k}$ and the normalized number of perceived moving objects ${\chi}_{\text{norm},k}$. The target output is the optimal priority weight $W_{\mathrm{target}}$. The loss function used for training is designed to minimize the discrepancy between the computed weights $w_{k}$ and the target weights $W_{\mathrm{target}}$, as described in Sec. IV-D. to compute priority weights based on normalized delay and the number of perceived objects ($\chi_{k}$):
@@
where $p_{k}$ denotes the computed priority score for camera $k$, and ${\Theta_{M}}$ represents the trainable parameters of MLP. The architecture of this MLP, featuring two layers, allows it to effectively model the interactions between delay and the number of perceived moving objects. Specifically, ${d}_{\text{norm},k}=\frac{{d}_{k}}{{d}_{\max}}$ and ${\chi}_{\text{norm},k}=\frac{{\chi}_{k}-{\chi}_{L}}{{{\chi}_{U}-{\chi}_{L}}}$, where ${\chi}_{k}$ represents the number of moving objects perceived by camera $k$, while ${\chi}_{U}$ and ${\chi}_{L}$ denote the upper and lower bounds of the number of moving objects that any edge camera should perceive, respectively.
@@ III-A Prioritized Information Bottleneck Analysis
In the context of information theory, the IB method seeks an optimal trade-off between the compression of an input variable $X$ and the preservation of relevant information about an output variable $Y$[49]. Throughout this paper, upper-case letters (e.g., $X$, $Y$, and $Z$) represent random variables, while lower-case letters (e.g., $x$, $y$, and $z$) denote their realizations. We formalize the input data from camera $k$ as $X^{(k)}$, and the target prediction as $Y^{(k)}$, corresponding to the population in the dataset $\mathcal{D}$. The goal is to encode $X^{(k)}$ into a meaningful and concise representation $Z^{(k)}$, which aligns with the hidden representation $z^{(k)}$ that captures task-relevant features of multi-view content for prediction tasks. The classical IB problem can be formulated as a constrained optimization task:
$\max_{\Theta}$ $\sum_{k=1}^{K}{I}\left(Z^{(k)};Y^{(k)}\right)$ (5)
@@
where $I(Z^{(k)},Y^{(k)})$ denotes the mutual information between two random variables $Z^{(k)}$ and $Y^{(k)}$. $\Theta$ represents the set of all learnable parameters in the PIB framework, including ${\Theta_{M}}$ and the variational approximation in the following section. The mutual information is essentially a measure of the amount of information obtained about one random variable through the other random variable. $I_{c}$ is the maximum permissible mutual information that $Z^{(k)}$ can contain about $X^{(k)}$. The objective is to ensure that $Z^{(k)}$ captures the most relevant information about $X^{(k)}$ for predicting $Y^{(k)}$ while remaining as concise as possible. By introducing a Lagrange multiplier [^2] All Lagrange multipliers $\lambda$ are the same, and we only use trainable weight parameters to dynamically balance between accuracy and communication bottleneck. $\lambda$, the problem is equivalently expressed as:
$\max_{\Theta}\quad R_{IB}=\sum_{k=1}^{K}\left[{I}\left(Z^{(k)};Y^{(k)}\right)-\lambda{I}\left(X^{(k)};Z^{(k)}\right)\right],$ (6)
where $R_{IB}$ represents the IB functional, balancing the compression of $X^{(k)}$ against the necessity of accurately predicting $Y^{(k)}$. Next, we extend the IB framework to a multi-camera setting by introducing priority weights to the mutual information terms, adapting the optimization problem as follows:
$\min_{\Theta}\quad\sum_{k=1}^{K}\left[-I_{w}\left(Z^{(k)};Y^{(k)}\right)+\lambda I_{w}\left(X^{(k)};Z^{(k)}\right)\right],$ (7)
@@
$\begin{cases}I_{w}\left(Z^{(k)};Y^{(k)}\right)=w_{k}\cdot I\left(Z^{(k)};Y^{(k)}\right),\\ I_{w}\left(X^{(k)};Z^{(k)}\right)=e^{w^{0}-w_{k}}\cdot I\left(X^{(k)};Z^{(k)}\right),\\ \end{cases}$ (8)
where the non-negative value $w^{0}$ represents the maximum allowable weight for $w_{k}$. The first term with linear weights $I_{w}\left(Z^{(k)};Y^{(k)}\right)=w_{k}I\left(Z^{(k)};Y^{(k)}\right)$ is the weighted mutual information between the compressed representation $Z^{(k)}$ from camera $k$ and the target $Y^{(k)}$. This term can also be used to capture the semantic compression in raw data. The linear weighting with $w_{k}$ ensures the influence of each camera is proportional to its priority weight. Higher $w_{k}$ values increase the weight given to $I\left(Z^{(k)};Y^{(k)}\right)$ in the objective function, emphasizing cameras that provide high-quality data for accurate target prediction.
@@ III-B Combinatorial Multiarmed Bandit (CMAB) Problem
where $\mathcal{C}_{a}$ represents the set of cameras already selected, and $M_{\mathcal{C}_{a}\cup\{k\}}$ represents the MODA score when the $k$-th camera is added to the set $\mathcal{K}$. To incorporate submodularity, the reward function needs to reflect the diminishing returns property. Therefore, we define the reward function $r_{\mathcal{K}}(t)$ as:
@@ IV-A Architecture Summary
In this subsection, we outline the workflow of our PIB framework, designed for collaborative edge video analytics. As depicted in Fig. 4, the process starts with each edge camera (denoted by $k$) capturing raw video data $X_{t}^{(k)}$ and extracting feature maps. These cameras utilize priority weights $w_{k}$ to optimize the balance between communication costs and perception accuracy, adapting to varying channel conditions. The extracted features are then compressed using entropy coding and sent as a bitstream to the edge server for further processing. On the server (see Fig. 4), the video features are reconstructed using the shared parameters such as weights $w_{k}$ and the variational model parameters $q(Z_{t}^{(k)}|Z_{t-1}^{(k)},...,Z_{t-\tau}^{(k)})$. The server integrates these multi-view features to estimate $Y_{t}$, such as pedestrian occupancy and object detection. This approach leverages historical frame correlations through a multi-frame correlation model to enhance prediction accuracy.

### v2

@@ Abstract
Collaborative perception systems leverage multiple edge devices, such surveillance cameras or autonomous cars, to enhance sensing quality and eliminate blind spots. Despite their advantages, challenges such as limited channel capacity and data redundancy impede their effectiveness. To address these issues, we introduce the Prioritized Information Bottleneck (PIB) framework for edge video analytics. This framework prioritizes the shared data based on the signal-to-noise ratio (SNR) and camera coverage of the region of interest (RoI), reducing spatial-temporal data redundancy to transmit only essential information. This strategy avoids the need for video reconstruction at edge servers and maintains low latency. It leverages a deterministic information bottleneck method to extract compact, relevant features, balancing informativeness and communication costs. For high-dimensional data, we apply variational approximations for practical optimization. To reduce communication costs in fluctuating connections, we propose a gate mechanism based on distributed online learning (DOL) to filter out less informative messages and efficiently select edge servers. Moreover, we establish the asymptotic optimality of DOL by proving the sublinearity of their regrets. To validate the effectiveness of the PIB framework, we conduct real-world experiments on three types of edge devices with varied computing capabilities. Compared to five coding methods for image and video compression, PIB improves mean object detection accuracy (MODA) while reducing 17.8% and reduces communication costs by 82.65% under poor channel conditions.
@@ I-A Background
Video analytics is rapidly transforming various sectors such as urban planning, retail analysis, and autonomous navigation by converting visual data streams into useful insights[2]. A large number of video cameras produce vast amounts of video data continuously and often require real-time video stream[3]. Numerous developing applications such as remote patient care[4], video games[5], UAV sensing[6, 7] and virtual and augmented reality depend on the efficient analysis of video data with minimal delay[8].
@@
The increasing number of smart devices requires a computational paradigm shift towards edge computing. This approach involves processing data closer to its source, resulting in several benefits compared to traditional cloud-based paradigms, particularly reduced latency and bandwidth costs. Even a delay as short as second can lead to disastrous consequences. For example, interactive applications such as online gaming and video conferencing require latencies below 100 ms to ensure real-time feedback and seamless user experience[9]. Similarly, VR/AR applications demand extremely low latencies, often less than 20 ms, to prevent motion sickness and maintain a high-quality immersive experience[10]. Utilizing remote cloud services for data processing can result in significant latency increases, often exceeding 100 ms[11]. Moreover, the importance of privacy, particularly in regions with strict data protection laws such as the General Data Protection Regulation (GDPR), makes edge computing even more attractive[12]. According to the Ponemon Institute, 60% of companies express apprehension toward cloud security and decide to manage their own data onsites in order to mitigate potential risks[13].
@@
A key aspect of video analytics, particularly for Bird’s Eye View (BEV) applications, is accurately capturing pedestrian occupancy across multiple camera views[14, 15]. BEV representations rely on precise spatial context to depict ground-level scenes in order to minimize occlusions, blind spots, and viewpoint discrepancies. Pedestrian occupancy data enables reliable identification and localization of individuals within a shared view, refining collaborative perception and enhancing tasks such as detection and prediction in complex, dynamic environments. However, the integration of edge devices into video analytics also brings in many significant challenges[16]. The computational demands of deep neural network (DNN) models, such as GoogLeNet[17], which requires about 1.5 billion operations per image classification, place a substantial burden on the limited processing capacities of edge devices[18]. Additionally, the outputs from high-resolution cameras increase the communication load. For example, a 4K video stream requires up to 18 Gbps of bandwidth to transmit raw video data, potentially overwhelming wireless networks[19]. Therefore, we need to explore efficient video coding for compressing streamed videos. As shown in Fig. 1(a), the traditional compression is to reconstruct streaming frame through efficient entropy coding and motion prediction. However, there are still extensive less informative data being processed, wasting communication bandwidth. For instance, if the tasks involve human recognition or positioning, the reconstructed background of each frame might not be useful for the application.
@@
In this paper, we aim to refine multi-camera video analytics by developing a strategy to prioritize wireless video transmissions. Our proposed Prioritized Information Bottleneck (PIB) strategy attempts to effectively leverage SNR and RoI to selectively transmit data features, significantly reducing computational load and data transmissions. Our method can decrease data transmissions by up to 82.65%, while simultaneously enhancing the mean object detection accuracy (MODA) compared to current state-of-the-art techniques. This approach not only compresses data but also intelligently selects data for processing to ensure only relevant information is transmitted, thus mitigating noise-induced inaccuracies in collaborative sensing scenarios. This innovation sets a new benchmark for efficient and accurate video analytics at the edge.
@@ I-C Our Contributions
  • Our extensive experimental evaluations across different real-world hardware platforms demonstrate that PIB significantly enhances mean object detection accuracy (MODA) and reduces communication costs. Compared to TOCOM-TEM, JPEG, H.264, H.265, and AV1, PIB improves MODA by 17.8% while reducing communication costs by 82.65% under poor channel conditions. Additionally, our method can reduce the standard deviation of streaming packet sizes by up to 9.43%, while simultaneously maintaining higher MODA, ensuring better transmission robustness under poor channel conditions.
@@ II System Model
As illustrated in Fig. 2, our system comprises a set of edge cameras, denoted as $\mathcal{K}=\{1,2,\ldots,K\}$, and a set of edge servers, denoted as $\mathcal{S}=\{1,2,\ldots,S\}$. Each camera has a specific Field of View (FoV), $\text{FoV}_{k}$, covering a subset of the monitored area. Our goal is to enable collaborative perception for pedestrian occupancy prediction under constrained channel capacity. The backbone of our system employs intermediate collaboration, where only encoded feature representations $Z$ are transmitted to a central aggregate server for decoding and inference. Other edge servers serve solely as relay nodes, as they do not perform decoding or final inference due to the network structure while does not support task offloading for these operations. The aggregate server is chosen to minimize overall system delay, while the relay-only role of other edge servers ensures streamlined data flow. The key notations are given in Table I for the ease of reading.
TABLE I: Key Notations
Notation Description
$C_{k,s}$ Capacity of the link between camera $k$ and edge server $s$
$d_{k,s}^{T}$ Transmission delay from camera $k$ to edge server $s$
$d_{k,s_{0}}^{I}$ Inference delay at the aggregate edge server $s_{0}$
$d_{r,s_{0}}^{R}$ Relay delay from edge server $r$ to aggregate server $s_{0}$
$p_{k}$, $w_{k}$ Priority score and weight for camera $k$
$X^{(k)}$ Input data from camera $k$
$Z^{(k)}$ Feature extracted from camera $k$
$Y$ Random variable of the output
$\mathcal{E}_{k,s}^{c\rightarrow e}$ Connection between camera $k$ and edge server $s$
$\mathcal{E}_{s}^{e\rightarrow e_{0}}$ Connection between edge server $s$ and aggregate server $s_{0}$
$\mathcal{K}$ Set of cameras selected at time $t$
$\Psi^{F}_{s}$ Computational cost at edge server $s$
$\Psi^{T}_{s}$ Computational cost of forwarding features
$\Psi^{R}_{k}$ Remaining computing capacity of edge server $k$
$\mathcal{T}_{k,s}^{c\rightarrow e}$ Latency from camera $k$ to edge server $s$
$\mathcal{T}_{s}^{e\rightarrow e}$ Latency between edge server $s$ and aggregate server $s_{0}$
@@ II-B Priority Weight Formulation
Dynamic priority weighting is essential for optimizing network resource allocation, as various data sources require different levels of attention. Inspired by our previous work[14], we employ a dual-layer Multilayer Perceptron (MLP) [^1] The MLP is trained in a supervised learning manner, where the input features are the normalized delay ${d}_{\text{norm},k}$ and the normalized number of perceived moving objects ${\chi}_{\text{norm},k}$. The target output is the optimal priority weight $W_{\mathrm{target}}$. The loss function used for training is designed to minimize the discrepancy between the computed weights $w_{k}$ and the target weights $W_{\mathrm{target}}$, as described in Sec. IV-D. to compute priority weights based on normalized delay and the number of perceived objects ($\chi_{k}$).
@@
where $p_{k}$ denotes the computed priority score for camera $k$, and ${\Theta_{M}}$ represents the trainable parameters of MLP. The architecture of this MLP, featuring two layers, allows it to effectively model the interactions between delay and the number of perceived moving objects. Specifically, ${d}_{\text{norm},k}=\frac{{d}_{k}}{{d}_{\max}}$ and ${\chi}_{\text{norm},k}=\frac{{\chi}_{k}-{\chi}_{L}}{{{\chi}_{U}-{\chi}_{L}}}$. To account for the dynamic nature of the system, we periodically update $d_{\max}$. Specifically, we define $d_{\max}$ as $d_{\max}=\max_{k}(d_{k})+\Delta$, where $d_{k}$ is the delay for camera $k$ and $\Delta$ is a predefined threshold. This dynamic computation ensures that the normalized delay $d_{\text{norm},k}=\frac{d_{k}}{d_{\max}}$ remains within a reasonable range, preventing resource overcommitment. ${\chi}_{k}$ represents the number of moving objects perceived by camera $k$, while ${\chi}_{U}$ and ${\chi}_{L}$ denote the upper and lower bounds of the number of moving objects that any edge camera should perceive, respectively.
@@ III-A Prioritized Information Bottleneck Analysis
In the context of information theory, the IB method seeks an optimal trade-off between the compression of an input variable $X$ and the preservation of relevant information about an output variable $Y$[51]. Throughout this paper, upper-case letters (e.g., $X$, $Y$, and $Z$) represent random variables, while lower-case letters (e.g., $x$, $y$, and $z$) denote their realizations. We formalize the input data from camera $k$ as $X^{(k)}$, and the target prediction as $Y$, corresponding to the population in the dataset $\mathcal{D}$. The goal is to encode $X^{(k)}$ into a meaningful and concise representation $Z^{(k)}$, which aligns with the hidden representation $z^{(k)}$ that captures task-relevant features of multi-view content for prediction tasks. The classical IB problem can be formulated as a constrained optimization task:
$\max_{\Theta}$ $\sum_{k=1}^{K}{I}\left(Z^{(k)};Y\right)$ (5)
@@
where $I(Z^{(k)},Y)$ denotes the mutual information between two random variables $Z^{(k)}$ and $Y$. $\Theta$ represents the set of all learnable parameters in the PIB framework, including ${\Theta_{M}}$ and the variational approximation in the following section. The mutual information is essentially a measure of the amount of information obtained about one random variable through the other random variable. $I_{c}$ is the maximum permissible mutual information that $Z^{(k)}$ can contain about $X^{(k)}$. The objective is to ensure that $Z^{(k)}$ captures the most relevant information about $X^{(k)}$ for predicting $Y$ while remaining as concise as possible. By introducing a Lagrange multiplier [^2] All Lagrange multipliers $\lambda$ are the same, and we only use trainable weight parameters to dynamically balance between accuracy and communication bottleneck. $\lambda$, the problem is equivalently expressed as:
$\max_{\Theta}\quad R_{IB}=\sum_{k=1}^{K}\left[{I}\left(Z^{(k)};Y\right)-\lambda{I}\left(X^{(k)};Z^{(k)}\right)\right],$ (6)
where $R_{IB}$ represents the IB functional, balancing the compression of $X^{(k)}$ against the necessity of accurately predicting $Y$. Next, we extend the IB framework to a multi-camera setting by introducing priority weights to the mutual information terms, adapting the optimization problem as follows:
$\min_{\Theta}\quad\sum_{k=1}^{K}\left[-I_{w}\left(Z^{(k)};Y\right)+\lambda I_{w}\left(X^{(k)};Z^{(k)}\right)\right],$ (7)
@@
$\begin{cases}I_{w}\left(Z^{(k)};Y\right)=w_{k}\cdot I\left(Z^{(k)};Y\right),\\ I_{w}\left(X^{(k)};Z^{(k)}\right)=e^{w^{0}-w_{k}}\cdot I\left(X^{(k)};Z^{(k)}\right),\\ \end{cases}$ (8)
where the non-negative value $w^{0}$ represents the maximum allowable weight for $w_{k}$. The first term with linear weights $I_{w}\left(Z^{(k)};Y\right)=w_{k}I\left(Z^{(k)};Y\right)$ is the weighted mutual information between the compressed representation $Z^{(k)}$ from camera $k$ and the target $Y$. This term can also be used to capture the semantic compression in raw data. The linear weighting with $w_{k}$ ensures the influence of each camera is proportional to its priority weight. Higher $w_{k}$ values increase the weight given to $I\left(Z^{(k)};Y\right)$ in the objective function, emphasizing cameras that provide high-quality data for accurate target prediction.
@@ III-B Combinatorial Multiarmed Bandit (CMAB) Problem
where $\mathcal{C}_{a}$ represents the set of cameras already selected, and $M_{\mathcal{C}_{a}\cup\{k\}}$ represents the MODA score when the $k$-th camera is added to the set $\mathcal{K}$. To incorporate submodularity [^5] The submodularity of $r_{a}(t)$ can be proven in Appendix C of [14]., the reward function needs to reflect the diminishing returns property. Therefore, we define the reward function $r_{\mathcal{K}}(t)$ as:
@@ IV-A Architecture Summary
In this subsection, we outline the workflow of our PIB framework, designed for collaborative edge video analytics. As depicted in Fig. 4, the process starts with each edge camera (denoted by $k$) capturing raw video data $X_{t}^{(k)}$ and extracting feature maps. These cameras utilize priority weights $w_{k}$ to optimize the balance between communication costs and perception accuracy, adapting to varying channel conditions. The extracted features are then compressed using entropy coding and sent as a bitstream to the edge server for further processing. On the server (see Fig. 4), the video features are reconstructed using the shared parameters such as weights $w_{k}$ and the variational model parameters $q(Z_{t}^{(k)}|Z_{t-1}^{(k)},...,Z_{t-\tau}^{(k)})$. The server integrates these multi-view features to estimate $Y_{t}$, such as pedestrian occupancy and object detection. This approach leverages historical frame correlations through a multi-frame correlation model to enhance prediction accuracy. The DNN architecture of the PIB framework is detailed in Appendix A.